A signal of frequency 10k Hz is being digitalized by an A/D converter. A possible
sampling time which can be used is
1
Expert's answer
2020-09-29T10:10:30-0400

Since a sample is a piece of data taken from whole data that is continuous in the time domain.

When the source generates an analog signal and if it is to be digitized having 1 s and 0 s, that is, High or Low, the signal must be time sampled. This sampling of the analog signal is called sampling.

That signal period is the sampling time, therefore

since, "v=\frac{1}{T},"where "v" is the frequency and is the period, "T" it is obtained that the sampling time is "T=\frac1v=\frac{1}{10^4}""=0,0001 seconds"